The coefficient of static and kinetic friction between a 3.0 kg box and a table are 0.40 and 0.30 respectively. What is the net force on the box when a 10.0 N horizontal force is applied to the box when at rest on a table?
1
Expert's answer
2020-01-03T09:18:34-0500

The static friction "F_s=\mu_s mg=0.40\times 3.0\times 9.8=11.8\:\rm N."

The kinetic friction "F_k=\mu_s mg=0.30\times 3.0\times 9.8=8.82\:\rm N."

Since the applied force is less than static friction, so the box will be at rest, and the net force is zero.